OpenStudy (anonymous):

Please Halp Meh! Which is the sum of 3/10 and 1/3?

OpenStudy (anonymous):

:P

OpenStudy (anonymous):

You have the make the fractions have a common denominator.

OpenStudy (anonymous):

3/10: multiply the numerator and the denominator by 3 = 9/30 1/3: multiply the numerator and denominator by 10 = 10/30

OpenStudy (anonymous):

9/30+10/30 = 19/30